OpenStudy (darkbluechocobo):

Apply the relationships between the quantities of frequency, wavelength, period, and wave speed to deterimine the period and wavelength of a wave with a speed of 15 m/s and a frequency of 300Hz. Show your reasoning.

OpenStudy (darkbluechocobo):

@Data_LG2

OpenStudy (shamim):

Given v=15m/s f=300hz

OpenStudy (shamim):

T=? Lamda=?

OpenStudy (shamim):

We know T=1/f v=f*lamda

OpenStudy (shamim):

Response plz for getting bettet help!!!!

OpenStudy (darkbluechocobo):

Sorry it was like 2am when i posted so I kinda just went to sleep

OpenStudy (darkbluechocobo):

T=0.00333333333 @shamim

OpenStudy (darkbluechocobo):

what is lamda?

OpenStudy (shamim):

T=1/f=1/300=?

OpenStudy (shamim):

v=f*lamda Lamda=v/f=15/300=?

OpenStudy (darkbluechocobo):

So T=1/300=0.0033 Lamda=15/300=0.05

OpenStudy (shamim):

Ya u r correct!!!!

OpenStudy (darkbluechocobo):

Thank you for the help :D!

OpenStudy (shamim):

U r most welcome!!!!!!"!